OpenStudy (anonymous):

If f(x) = 5x2 − 4 and g(x) = x − 2, what is (f o g)(-3)

OpenStudy (bahrom7893):

well: (f o g)(-3) = f(g(-3)) g(-3) = -3 - 2 = -5 Thus f(g(-3)) = f(-5) = 5*(-5)^2 - 4 = 125 - 4 = 121